Hi
Betty has 1/3 as many pins as susan.when susan gave away 55 pins, the number of pins Betty has is twice the number susan has left..how many pins does susan have at first.
thanks

Answer by jorel555(1290)

Let s be susan, b be betty. Then:
s=3b
s-55=b/2. So:
3b-55=b/2
6b-110=b
5b=110
b=22
s=66
Betty has 22 pins to start, Susan has 66!!!!!!!!!!!!!!